OpenStudy (anonymous):

-4p-10(2p-3)=22-7(1-2p)

OpenStudy (anonymous):

Shenita, I can help with this..... ok, first distribute the minus signs in front of the 10 and 7 when you simplify the equation....you'll get -4p-20p+30=22-7+14p See so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then continue simplifying like terms on each side....remember that -4p-20p=-24p and you'll get -24p+30=15+14p With me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now start to do your inverse operations...I usually teach my students to do the variables first, but it doesn't really matter....and I start with the SMALLER of the variables....that will help avoid some negatives ....So I would ADD 24p to both sides (I add them underneath each side....to get 30= 15 +38p...see this? Write it with adding the 24p underneath each variable term

OpenStudy (anonymous):

See how -24p is smaller than 14p??

OpenStudy (anonymous):

yes negative

OpenStudy (anonymous):

Now do the inverse of a positive 15 to both sides which is to subtract 15 (underneath again)...and you'll get 15=38p

OpenStudy (anonymous):

just checking my math...are you getting otjer answers with fractions?

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

when you divide both sides by 38 (the last step) you'll get 15/38=p

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which doesn't reduce....is this part of a homework assignment? Some books like to have fractional answers....some stay away from them...

OpenStudy (anonymous):

yes she wants it to stay as fractions

OpenStudy (anonymous):

yeah!! love that teacher!! PLEASE work with fractions as often as possible.....instead of making everything more difficult turning them into decimals!!!

OpenStudy (anonymous):

yes thanx laurie

OpenStudy (anonymous):

Answers actually make more sense in fraction form. :-) And you're welcome.